Can someone please help me with this problem I have been working on it all day.
Multiply
(x-2)(3x-6)(x+3)

Found 2 solutions by fabioGR, vasu2qute:
Answer by fabioGR(2)

(3x^2-6x-6x+12)(x+3)
(3x^2-12x+12)(x+3)
3x^3-12x^2+12x+9x^2-36x+36
3x^3-3x^2-24x+36

Answer by vasu2qute(17)

(x-2)(3x-6)(x+3)=
(x-2)[3(x-2)](x+3)=3[(x-2)^2](x+3)= 3(x^2-4x+4)(x+3)=
=3(x^3-3x^2-8x+12)